OpenStudy (anonymous):

Convert theta = 3pi/4 to cartesian form

OpenStudy (anonymous):

X=RCOS(THETA),Y=RSIN(THETA), X^2+Y^2=R^2

OpenStudy (anonymous):

is an example from polar to cartesian form

OpenStudy (anonymous):

which should give youthe hint

OpenStudy (anonymous):

Yes but how do I convert theta.? I have do I use inverse sin or cos?

OpenStudy (anonymous):

you can find tan theta and use inverse trig formula

OpenStudy (anonymous):

cos 3pi/4= -1\sqrt(2), x=-R/sqrt(2)

OpenStudy (anonymous):

dont you have value of R

OpenStudy (anonymous):

Ah, that's what was confusing me. I not having r that is

OpenStudy (anonymous):

if you do not have value of R, let it be as a variable

OpenStudy (anonymous):

it is not wise to assume radius as unity

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

yw